OpenStudy (anonymous):

Find the cube roots of 27(cos 279° + i sin 279°).

OpenStudy (anonymous):

cube root for the magnitude, divide by 3 for the angle...

OpenStudy (jdoe0001):

\(\large \bf { \textit{De Moivre's theorem}\\ Z^{\color{blue}{ n}}=[r[cos(\theta)+i\ sin(\theta)]]^{\color{blue}{ n}}\iff r^{\color{blue}{ n}}[cos({\color{blue}{ n}}\theta)+i\ sin({\color{blue}{ n}}\theta)] \\ \quad \\ \quad \\ thus\qquad \sqrt[{\color{blue}{ 3}}]{Z}\implies Z^{{\color{blue}{ \frac{1}{3}}}}\implies r^{\color{blue}{ \frac{1}{3}}}[cos({\color{blue}{ \frac{1}{3}}}\theta)+i\ sin({\color{blue}{ \frac{1}{3}}}\theta)] }\)

OpenStudy (dumbcow):

to get the other 2 roots, add multiples of 2pi to angle before dividing by 3 \[\cos (\frac{279 + 360k}{3}) + i \sin (\frac{279+360k}{3})\] k = 0,1,2